# Collaborating with AI Agents: Field Experiments on Teamwork, Productivity, and Performance

### v1

@@ Abstract
To uncover how AI agents change productivity, performance, and work processes, we introduce MindMeld: an experimentation platform enabling humans and AI agents to collaborate in integrative workspaces. In a large-scale marketing experiment on the platform, 2310 participants were randomly assigned to human-human and human-AI teams, with randomized AI personality traits. The teams exchanged 183,691 messages, and created 63,656 image edits, 1,960,095 ad copy edits, and 10,375 AI-generated images while producing 11,138 ads for a large think tank. Analysis of fine-grained communication, collaboration, and workflow logs revealed that collaborating with AI agents increased communication by 137% and allowed humans to focus 23% more on text and image content generation messaging and 20% less on direct text editing. Humans on Human-AI teams sent 23% fewer social messages, creating 60% greater productivity per worker and higher-quality ad copy. In contrast, human-human teams produced higher-quality images, suggesting that AI agents require fine-tuning for multimodal workflows. AI personality prompt randomization revealed that AI traits can complement human personalities to enhance collaboration. For example, conscientious humans paired with open AI agents improved image quality, while extroverted humans paired with conscientious AI agents reduced the quality of text, images, and clicks. In field tests of ad campaigns with ~5M impressions, ads with higher image quality produced by human collaborations and higher text quality produced by AI collaborations performed significantly better on click-through rate and cost per click metrics. Overall, ads created by human-AI teams performed similarly to those created by human-human teams. Together, these results suggest AI agents can improve teamwork and productivity, especially when tuned to complement human traits.
@@ 1 Introduction
Generative artificial intelligence (AI) tools have garnered attention for their potential to improve productivity and performance. Not only do they have far-reaching impacts across industries (Eloundou et al., 2024, Bick et al., 2024), but they also increase measurable productivity statistics, like output per unit time, as well as creativity and measurable quality. Researchers have found positive productivity effects of AI tools, with evidence of heterogeneity across tasks and baseline performance. For example, large language models (LLMs) decreased the average time taken for mid-level professional writing tasks by 40% and increased quality by 18% (Noy and Zhang, 2023). For job seekers, AI assistance with resumes increased job hiring by an average of 8% (Wiles et al., 2023), and for customer support workers, AI assistance increased productivity by an average of 14% (Brynjolfsson et al., 2023). Moreover, productivity gains were greater for lower-skilled workers (Noy and Zhang, 2023, Brynjolfsson et al., 2023, Choi and Schwarcz, 2023) and varied across task domains (Dell’Acqua et al., 2023). Evidence from an online labor market suggests there has already been a reduction in demand for freelance knowledge work with the advent of generative pre-trained transformer (GPT) models (Hui et al., 2023). In a study reviewing over 5,000 papers, Vaccaro et al. (2024) shows that human-AI groups outperform humans alone in 85% of the studies.
@@
This study addresses three critical gaps in understanding AI’s impact on productivity and collaboration: 1) the lack of understanding of the role of AI agents, 2) the dearth of evidence concerning the mechanisms behind productivity improvements and how work processes change when humans collaborate with AI, and 3) whether and when model prompt engineering can improve the complementarity between AI agents and human labor. First, while the current literature, such as Dell’Acqua et al. (2023) and Chen and Chan (2024), reveal the productivity effects of AI by randomizing access to LLM chatbots, they are not multimodal, do not include context, do not allow the chatbots to take independent actions or use APIs to call outside of the platform, and do not provide a collaborative workspace where machines and humans can jointly manipulate output artifacts in real-time. These innovations are meaningful because AI agents today have all these features, yet the existing scientific literature studies none of them. Moreover, while Liu et al. (2023) explores levels of AI proactivity, which is key to understanding the effects of AI agents, it is a demonstration paper without a randomized controlled trial (RCT) measuring productivity effects. Similarly, Chen and Chan (2024) examines artificial workflows that do not reflect real-world settings.
@@
Second, we lack insight into how work processes and collaboration patterns change when humans collaborate with AI agents compared to human-only teams and how these changes affect productivity and performance. Existing research has focused on the productivity effects of GPT chatbots (Noy and Zhang, 2023, Dell’Acqua et al., 2023). Separately, others investigate how AI changes people’s perceptions, beliefs, and behaviors (Tey et al., 2024, Costello et al., 2024). However, it is unclear how these interactions evolve in real-time collaboration settings, especially in environments where AI agents can take autonomous actions, adapt dynamically to human inputs, and participate in tasks requiring creativity and coordination. Current off-the-shelf experimental platforms and studies do not provide collaborative workspaces where researchers can precisely record and measure the collaboration itself: e.g., transcripts of messages between machines and humans, logs of edits to output artifacts, and API (application programming interface) calls to outside agents. In contrast, current AI applications, such as Notion AI, v0.dev, OpenAI Canvas, Cursor, and GitHub Copilot, already integrate AI agents in collaborative workspaces in addition to chat interfaces.
Third, prompts are a critical factor in LLM performance and have been studied extensively (Schulhoff et al., 2024). Indeed, the seminal paper introducing the GPT-3 model highlights that much of its power lies in its ability to learn and adapt to the context provided within the prompt, allowing it to generate coherent text without task-specific fine-tuning (Brown et al., 2020). Since then, other prompting methods, such as chain-of-thought prompting (Wei et al., 2023), have been shown to significantly improve task performance. However, little is known about how different prompting strategies interact with humans in collaborative workflows. In the human-robot interaction literature, studies have shown that factors like a robot’s social communication and appearance influence the effectiveness of human-robot teams (Pamela J. Hinds and Jones, 2004, Jung et al., 2013, Jung and Hinds, 2018). For GPT-based agents, prompt engineering can be similarly critical; thus, it is essential to understand prompt effects on human-agent collaboration. Rigorous randomized controlled trials (RCTs), such as Jakesch et al. (2023), are needed to systematically evaluate the effects of prompts on human-AI collaboration and task outcomes.
@@
To address these gaps, we developed MindMeld, a novel experimental platform designed to study human-AI collaboration in real-world, extensible tasks. MindMeld introduces several key innovations. It enables real-time collaboration between humans and AI agents, allowing participants to manipulate text, images, and workflows collaboratively in a chat-enabled workspace that mirrors existing online collaboration workflows. The platform supports randomized pairings of humans and AI (i.e., Human-Human or Human-AI pairs) and allows for randomization of prompts. Critically, the AI can perform an equivalent set of actions that humans can on the platform. These include sending chat messages, editing the copy, selecting images, and generating new images using an external call to Dall-E 3. Thus, MindMeld captures every time-stamped keystroke, message, edit, API call, and intermediate output, providing a rich dataset that allows the detailed reconstruction of collaboration workflows. This represents a fundamental departure from the existing literature, which enables RCT-based evaluation of chatbots and co-pilots (LLMs) but does not enable randomized experiments analyzing human collaboration with AI agents.
@@
We conducted a large-scale study using MindMeld to examine the effects of human-AI collaboration on ad design—a task requiring creativity, iteration, and precision. A total of 2,310 participants, representative of the U.S. population, were recruited through Prolific and randomly paired into human-human or human-AI teams. As research on AI agents at this level of capability is relatively new, our randomization of human-AI teams serves as a bundle intervention, since AI agents exhibit many behaviors different from those of humans, with which we can use for further hypothesis generation. AI agents, built on GPT-4o, were further randomized by prompts designed to induce high or low levels of Big Five personality traits, enabling us to investigate how personality alignment influences collaboration, productivity, and quality. Teams worked collaboratively to create marketing campaigns for a real organization’s year-end annual report, including generating ad images, writing text copy, and editing headlines. This process was fully recorded, resulting in 11,138 ads created, 183,691 messages exchanged, 1,960,095 text edits, 63,656 image edits, and 10,375 AI-generated images, offering an unprecedented level of detail to analyze teamwork dynamics.
@@
Once the lab portion of the experiment was completed, we turned to the field evaluation. We first obtained human and AI quality ratings of the ads, including the overall quality of the ads, the quality of the images, the ad copy, the headlines, and the self-reported or AI-evaluated likelihood of consumer engagement with the ads (click-through rates). The ad copy was evaluated by human and AI raters on clarity, conciseness, grammar mistakes, persuasiveness, relevance to task, tone, and voice, the strength of the call to action, and the degree to which the ad was “attention-grabbing.” We then ran the ads online, on X, generating over 4.9 million impressions, and evaluated click-through rates, cost-per-click, and view-through rates on the annual report, using DocSend, which allows us to see how much of the report consumers read, page by page, after clicking through on the ads.
Building on the capabilities of MindMeld, this study investigates how AI agents reshape collaboration dynamics, productivity, and performance compared to human-only teams. While prior work has shown that AI tools enhance productivity and reduce task completion times (Noy and Zhang, 2023, Brynjolfsson et al., 2023), these studies often treat AI as a passive tool rather than an active collaborator. As AI agents become integral to workflows, researchers are beginning to explore their roles as counterparts in collaborative systems rather than mere tools or mediums, emphasizing the importance of trust, transparency, and integration in human-AI partnerships (Makarius et al., 2020, Anthony et al., 2023, Collins et al., 2024). We hypothesize that AI agents, by reducing social coordination costs and enabling participants to focus more on content generation, will enhance individual productivity and lead to distinct communication patterns compared to human-human teams. Additionally, we investigate how the alignment between human personality traits and AI-induced traits influences collaboration outcomes. By randomizing AI prompts to induce high or low levels of Big Five personality traits, we examine how personality compatibility affects productivity, creativity, and collaboration quality. For instance, we hypothesize that high-openness pairings between humans and AI agents will lead to greater idea generation. These questions address critical gaps in understanding how AI agents interact with humans in multimodal workflows and how their design can be optimized for collaborative success.
@@
Our findings reveal profound differences in how AI reshapes teamwork and productivity. We found that collaborating with AI agents significantly reshaped teamwork dynamics and enhanced productivity and performance compared to the human-only teams. In mining the rich collaboration data to understand how AI changes teamwork, we found Human-AI teams sent 45% more messages than Human-Human teams, revealing increased communication. The results were also consistent at the individual level: humans in Human-AI teams messaged more than those in Human-Human teams. Furthermore, both AI and humans in Human-AI teams sent more content- and process-oriented messages about their collaboration than those in Human-Human teams, especially messages containing suggestions, instructions, prioritization, judgment, and planning. Conversely, humans in Human-Human teams sent more messages that were social and emotional—including more messages that expressed rapport building, self-assessment, and concern—than humans in Human-AI teams. These results begin to unpack how work changes when humans collaborate with AI agents compared to when they collaborate with humans. We also tracked the work done on the collaborative workspace where humans and agents could both collaboratively edit the ad copy as well as generate images using external APIs (i.e., Dall-E 3). We found that Human-AI teams made 84% fewer edits to the copy due to LLM’s proficiency in writing high-quality ad copy. In contrast, LLMs were worse at predicting image quality, thus disadvantaging human-AI teams in the visual dimensions of ad creation and leading to human-AI teams producing lower-rated images. These differences in collaboration outcomes indicate a shift in workload that allowed human participants in Human-AI teams to engage more in content generation with less social coordination effort, including such coordinating activities like rapport building, than human participants in Human-Human teams.
The implications of these collaboration differences are then evident in the productivity and performance differences. Human-AI teams produced as many ads as Human-Human teams and 70% more ads on the individual level. First, we found in AI evaluations by gpt-4o-mini that ads created in Human-AI teams scored higher on the quality of the text but equally on image quality. Interestingly, human evaluations (using a separate set of 1,195 Prolific workers) also scored text quality higher but lower on image quality. Both human and AI evaluations are tied to the estimated likelihood of click-through rates. In addition to randomizing whether humans were paired with AI, we also orthogonally randomized a set of agent prompts designed to induce AI with either high or low levels of each of the Big Five personality traits. Through prompt randomization, we found heterogeneous prompt effects: a high or low personality prompt had differential effects on our outcomes depending on the personality of the human participant.
Overall, these findings suggest that AI’s involvement in collaborative settings can drive productivity and performance by enabling participants to focus more on the content, possibly by reducing the social coordination costs of collaboration. Importantly, these effects vary across the personality of human collaborators, which means AI agents can be tuned to “fit” the personality of human collaborators to improve productivity and performance. By combining the precise manipulability of lab experiments with the rigor and real-world measurability of field studies, this study provides novel insights into the dynamics and potential of collaborating with AI agents.
2 Methods
@@
Our study was preregistered and deemed minimal risk and exempt by the MIT Committee on the Use of Humans as Experimental Subjects (protocol E-5927). [^1] See osf.io/jfzha and osf.io/95dhu. Any non-pre-registered analyses are labeled post hoc.
@@ 2.1 Procedure
As soon as a participant was redirected to our platform from Prolific, they were randomized into either the Human-Human or the Human-AI condition (Figure 1A). In the Human-Human condition, a participant joined a queue until another participant also joined the queue, at which point they were paired with each other. In the Human-AI condition, a participant joined a simulated queue, in which they waited for a random amount of time between 1 and 5 seconds, after which they were paired with an AI agent. We do not reveal whether or not the partner is a human or an AI until the post-task survey.
Figure 1: Overview of methods. (A) Participants are randomized into collaborating with another participant or an AI agent. (B) AI agents are assigned a personality profile based on Big Five traits, with each trait randomly set to either a high or low level. (C) Participants collaborate with another participant or an AI agent to produce ads in a real-time collaborative workspace.
Personality prompt randomization.
In addition to the context of the task and chat, we prompt the model with personality prompts for the Big Five personality traits (McCrae and John, 1992). For each trait, we randomize whether the personality is high ($p=0.5$) or low ($p=0.5$; Figure 1B). To induce high or low personality traits, we use prompts generated by Jiang et al. (2023) using P2 prompting, which uses a chain of prompts to generate a detailed description of individuals with the traits. The prompts are available on their GitHub repository.
@@
After participants were paired with each other, they answered a 10-item survey, each on a 7-point Likert scale, to measure their Big Five personality traits (Rammstedt and John, 2007).
@@
Once the participants were paired and completed the pre-task survey, they entered the collaborative workspace and chat interface in which they could message each other and edit and submit ads (Figure 1C). The edits were synchronized and messages were transmitted between the participants in real-time using websockets (i.e., tiptap.dev, pusher.com), as in commercial collaboration tools (e.g., Google Docs) and chat applications (e.g., Slack). The participants had 40 minutes to submit and could submit zero to as many ads as they could produce. At the end of the 40 minutes, the participants were automatically redirected to the post-task survey.
@@
After the completion of the ad creation task, the participants answered a 35-item teamwork quality survey (Hoegl and Gemuenden, 2001). The original survey consisted of 38 items, each on a 7-point Likert scale; however, we removed 3 items from the communication facet that did not apply to a single-session online collaboration task. The excluded items include the following: “The team members communicated often in spontaneous meetings, phone conversations, etc.”, “The team members communicated mostly directly and personally with each other.”, and “There were mediators through whom much communication was conducted.”
@@
In addition, we asked four questions regarding the perception of AI, all on a 7-point Likert scale. They include two on their experience using AI (i.e., “I have used artificial intelligence (AI) chatbots before (e.g., ChatGPT, Bard).” and “I had a positive experience using AI chatbots.”). The last two questions include one on their perception of their partner as an AI (i.e., “I believe my partner was an AI during the task.”) and the other in which we revealed the identity of their partner and whether that changed the perception of the quality of the collaboration (i.e., “Your partner was [an AI assistant/a human]. Knowing this, to what extent has your perception of the quality of your collaboration changed?”). Only 1964 participants completed the post-task survey.
2.2 The MindMeld platform
To study the impact of collaborating with AI agents, we developed an online platform which we call MindMeld (Figure 2). Once a participant enters the platform, the participant is randomized into a queue for either a Human-Human or Human-AI condition. In the Human-Human condition, the participant collaborates with another participant; in the Human-AI condition, the participant collaborates with an AI agent. As far as we know, this randomization is a unique feature of our platform.
Figure 2: The MindMeld platform. On the left is the task panel, and on the right is the chat panel. In the Human-Human condition, chat messages and edits on the task panel, including text edits, image selections, and AI image generations, are synchronized in real-time. In the Human-AI condition, the participant chats with an AI agent with full context of the user interface (UI; see Section 2.3), and the AI can edit text, select images, and generate AI images.
On the left side of MindMeld (Figure 2) is the task panel where participants can create, edit, and submit the ads. The platform includes a selection of images participants can choose from, and the participants can also access the Dall-E 3 application programming interface (API) to generate new images. The copy includes a headline, primary text, and a description. All of the edits—including the image selection and image generation as well as the copy edits—are synchronized in real-time across the participants. On the right is the chat panel in which participants can chat in real-time with either a participant or an AI agent. The participants can submit the ads, and the interface will reset so they can submit the next ad. To the best of our knowledge, our platform is the first to simulate real-time collaboration between human-human or human-AI pairs with real-time chat and editing on synchronized text and image interfaces.
2.3 AI agent
Context.
The platform uses OpenAI’s multimodal gpt-4o model (specifically gpt-4o-2024-08
-06
), which processes text and images in a single neural network. To give the AI full context of the task and collaboration, each API call to gpt-4o is prompted with the information on the screen so it has the full context of the collaboration. In the prompt, we include the following: the same text of the task given to participants, previous submissions, personality prompts (see Section 2.1), current copy, elapsed time, history of its own actions (see Section 2.3), history of its chain-of-thought (see Section 2.3), chat history, and general instructions. Moreover, a screenshot is taken of the image after every change and included as input so the AI can observe the image. The prompt template is shown in Appendix B.
Actions.
To ensure that the Human-Human and Human-AI conditions are comparable, the AI agent can take the same actions a human participant can except for submitting ads. The actions include the following: sending messages, editing any of the ad copy (i.e., headline, primary text, and description), selecting images, generating images using the Dall-E 3 interface, and waiting (i.e., not taking any action). The agent is prompted every 10 seconds whether to engage in action.
Chain-of-thought.
To ensure the agent is taking actions appropriately, we use chain-of-thought prompting (Wei et al., 2023). [^2] We used OpenAI’s structured output feature for chain-of-thought prompting. See their documentation. Specifically, we prompt the model with questions to reflect on the state of the collaboration. These questions were necessary to avoid undesirable behaviors of the model, such as repeatedly sending the same message, and were determined through trial and error. The chain-of-thought prompts are shown in Appendix B.

### v2

# Collaborating with AI Agents: A Field Experiment on Teamwork, Productivity, and Performance

@@ Abstract
To uncover how AI agents change productivity, performance, and work processes, we introduce Pairit—an experimentation platform enabling humans and AI agents to collaborate in integrative workspaces. In a large-scale marketing experiment on the platform, 2310 participants were randomly assigned to human-human and human-AI teams. The teams exchanged 183,691 messages and created 63,656 image edits, 1,960,095 ad copy edits, and 10,375 AI-generated images while producing 11,138 ads for a large think tank. Analysis of fine-grained communication, collaboration, and workflow logs revealed that collaborating with AI agents increased communication by 63% and allowed humans to engage in 71% less direct text editing. While human-AI teams engaged in 18% more process and content communication, human-human teams engaged in 29% more social and emotional communication. Humans in human-AI teams experienced 73% greater productivity per worker and produced higher-quality ad copy, while human-human teams produced higher-quality images, suggesting AI agents require fine-tuning for multimodal workflows. Field tests of the ad campaigns accumulated $\sim 5\text{M}$ ad impressions and revealed that ads with higher image quality (produced by human-human collaborations) and higher text quality (produced by human-AI collaborations) performed significantly better on click-through rates, view through rates, and cost per click metrics. Together, these results suggest that human collaboration with AI agents significantly reshapes communication patterns and work processes and increases productivity, while improving some dimensions of output quality and deteriorating others. We hope the release of the extensible Pairit platform will accelerate RCTs of human-AI collaboration across a variety of work tasks and contexts.
@@ 1 Introduction
Artificial intelligence (AI) tools have garnered attention for their potential to improve productivity and performance (Eloundou et al., 2024, Bick et al., 2024). For example, large language models (LLMs) decreased the average time taken for mid-level professional writing tasks by 40% and increased output quality by 18% (Noy and Zhang, 2023). For job seekers, AI assistance with resumes increased job hiring by an average of 8% (Wiles et al., 2023), and for customer support workers, AI assistance increased productivity by an average of 14% (Brynjolfsson et al., 2023). Moreover, productivity gains were greater for lower-skilled workers (Noy and Zhang, 2023, Brynjolfsson et al., 2023, Choi and Schwarcz, 2023) and varied across task domains (Dell’Acqua et al., 2023). Evidence from an online labor market suggests there has already been a reduction in demand for freelance knowledge work with the advent of generative pre-trained transformer (GPT) models (Hui et al., 2023). In a study reviewing over 106 papers, Vaccaro et al. (2024) showed that human-AI groups outperformed humans alone in 85% of the studies. Our work aims to address two critical gaps in this burgeoning literature on AI’s impact on collaboration and productivity: Understanding 1) the role of multimodal AI agents (as distinct from LLM-enabled chatbots) and 2) the work process changes created by human-AI agent collaboration.
@@
While studies like Liu et al. (2023) explore levels of proactivity in AI agents, they typically do so without randomized controlled trials (RCTs) measuring productivity effects. The studies that use RCTs to estimate productivity effects tend to randomize access to LLM chatbots (e.g. Dell’Acqua et al. (2023) and Chen and Chan (2024)), which are not typically multimodal, do not include context, do not allow the chatbots to take independent actions or use APIs to call outside of the work environment, and do not provide a collaborative workspace where machines and humans can jointly manipulate output artifacts in real time. These innovations are meaningful because AI agents today possess all these features, yet the existing scientific literature studies none of them.
@@
Furthermore, we currently lack fine-grained task level insight into how human-AI collaboration changes work processes and communication patterns and how these changes affect productivity and performance. The vast majority of existing research focuses on the productivity effects of GPT chatbots on individual workers (Noy and Zhang, 2023, Dell’Acqua et al., 2023). Others investigate how AI changes people’s perceptions, beliefs, and behaviors (Tey et al., 2024, Costello et al., 2024). However, it is unclear how these interactions evolve in real-time collaborations, especially in environments where AI agents can take autonomous actions, adapt dynamically to human inputs, and participate in tasks requiring creativity and coordination. Current off-the-shelf experimental platforms and studies do not provide collaborative workspaces where researchers can precisely record and measure the collaboration itself: e.g., transcripts of messages between machines and humans, logs of edits to output artifacts, and API (application programming interface) calls to outside agents. In contrast, current AI applications, such as Notion AI, v0.dev, OpenAI Canvas, Cursor, and GitHub Copilot, already integrate AI agents into such collaborative workspaces and interfaces.
@@
To address these gaps, we developed Pairit, a novel experimental platform designed to study human-AI collaboration in real-world, extensible tasks. Pairit introduces several key innovations. It enables real-time collaboration between humans and AI agents, allowing participants to manipulate text, images, and workflows collaboratively in chat-enabled workspaces that mirror existing online AI collaboration work processes. The platform supports randomized pairings of humans and AI (i.e., human-human or human-AI teams) and allows for randomization of prompts and model fine-tuning. Critically, the AI can perform the complete set of equivalent actions that humans can perform in the collaborative workspace. These include sending chat messages, writing ad copy, editing ad copy, writing calls to action, editing calls to action, scrolling through images, editing images, selecting images, and generating new images using an external call to Dall-E 3. Moreover, Pairit captures every time-stamped keystroke, message, edit, swipe, scroll, selection, API call, and intermediate output, providing a rich dataset that allows for the detailed reconstruction of collaboration workflows. This represents a fundamental departure from the existing literature, which enables RCT-based evaluation of the outputs and productivity implications of human collaboration with LLM-based chatbots and co-pilots, but does not enable randomized experiments analyzing the task-level productivity and work process changes created in human collaborations with fully functioning, multimodal AI agents.
@@
To test the Pairit platform and to study the work process, productivity, and performance implications of human collaboration with fully functional AI agents, we conducted a large-scale randomized study of the effects of human-AI collaboration on advertising design and creation—a task requiring creativity, iteration, and precision. A total of 2,310 participants, representative of the U.S. population, were recruited through Prolific and randomly paired into human-human or human-AI teams. Teams worked collaboratively to create marketing campaigns for a think tank’s year-end annual report, including generating and selecting ad images and writing ad copy and calls to action. This process was fully recorded, resulting in a dataset including 11,138 ads, 183,691 messages, 1,960,095 text edits, 63,656 image edits, and 10,375 AI-generated images, offering an unprecedented level of detail with which to understand work processes, teamwork dynamics, communication, productivity and output quality.
@@
Once the lab portion of the experiment was completed, we conducted a field experiment on the ads produced by human-human and human-AI teams. We obtained human and AI quality ratings of the ads, including the quality of the ad copy and images, and the human- and AI-evaluated likelihood of consumer engagement with the ads (measured by click-through rates). We then ran the ads a real online display ad ecosystem, generating over 4.9 million impressions, and evaluated click-through rate, cost-per-click, view-through rate and view through duration metrics on the annual report using the platform’s ads API and DocSend’s view metrics, which allow us to record how much of the report consumers read, page by page, after clicking through on the ads.
In mining the rich collaboration data to understand how AI changes team communication, we found individuals in human-AI teams sent 63% more messages than those in human-human teams and that results were consistent at the team level. Furthermore, both the humans and the AI in human-AI teams sent 18% more content- and process-oriented messages than those in human-human teams, especially messages containing suggestions, instructions, prioritization, and planning. Conversely, humans in human-human teams sent 29% more social and emotional messages, including more messages that expressed rapport building, self-assessment, and concern, than humans in human-AI teams. These results demonstrate that human-human teams spent considerable time and effort on social and emotional calibration and regulation, while human-AI teams were task- and process-oriented.
@@
We also tracked work process changes in the collaborative workspace where humans and agents could both edit the ad copy as well as generate images using external APIs (i.e., Dall-E 3). We found that human-AI teams made 71% fewer edits to the copy due to AI’s proficiency in writing high-quality ad copy. In contrast, the AI was worse at predicting image quality, thus disadvantaging human-AI teams in the visual dimensions of ad creation and leading to human-AI teams creating lower-rated images. These differences in collaboration outcomes indicate a shift in workload that allowed human participants in human-AI teams to engage more in content generation with less social coordination, like rapport building, than humans in human-human teams.
These communication, collaboration, and work process changes then translated into measurable differences in productivity and performance. Humans in human-AI teams experienced 73% greater productivity per worker and human evaluations scored the ads created by human-AI teams higher on text quality but lower on image quality. AI evaluations scored the ads created by human-AI teams higher on text quality and equally on image quality. Field tests of the ad campaigns revealed that ads with higher image quality (produced more often by human collaborations) and higher text quality (produced more often by human-AI collaborations) performed significantly better on click-through rates, view through duration and cost per click metrics. As human-AI teams produced higher quality text and lower quality images, human-AI ads performed similarly to those created by human-human teams overall. Together, these results suggest that human collaboration with AI agents significantly reshapes communication patterns and work processes and increases productivity, while improving some dimensions of output quality and deteriorating others.
Although prior studies have shown that AI tools can improve productivity and reduce task completion times (Noy and Zhang, 2023, Brynjolfsson et al., 2023), AI is often treated as a passive tool rather than as an active collaborator. As AI agents become integral to modern workflows, researchers are beginning to explore their role as work collaborators, rather than mere tools, emphasizing the importance of trust, transparency, and integration in human-AI partnerships (Makarius et al., 2020, Anthony et al., 2023, Collins et al., 2024). Our work contributes to this emerging literature by presenting the first task-level randomized experiment measuring the work process, productivity and performance implications of human-AI collaboration with fully functional, multimodal AI agents. We hope the release of the extensible Pairit platform will accelerate the deployment of RCTs of human-AI collaboration across a variety of work tasks and contexts in future research.
2 Methodology
@@
Our study was preregistered and approved by the MIT Committee on the Use of Humans as Experimental Subjects. [^1] See osf.io/jfzha and osf.io/95dhu. Any non-pre-registered analyses are labeled post hoc.
2.1 The Pairit platform
Once a participant enters the Pairit (Figure 1) platform, they are randomized into a queue for either the human-human or the human-AI condition. In the human-human condition, the participant collaborates with another human participant. In the human-AI condition, the participant collaborates with an AI agent. In addition to team-level randomization, the AI models are also configured to accept prompt and fine-tuning randomization. Pairit’s left panel is the task panel in which participants can participate in the task work process (in the context of this experiment to create, edit, and submit ads). The platform includes a carousel of images participants can choose from, and the participants (as well as the AI) can also access the Dall-E 3 application programming interface (API) to generate new images. The ad copy creation and editing panel includes a headline, primary text, and a description. All of the edits—including the image selection and image generation, as well as the ad copy edits—are synchronized in real-time across the participants, whether they are human or AI. On the right is the chat panel in which participants can chat in real time with either another human participant or an AI agent. The participants can submit the ads and the interface will reset so that they can work on and then submit the next ad. To the best of our knowledge, Pairit is the first platform to faciliate real-time collaboration between human-human or human-AI teams with team- and model-level randomization and real-time chat and editing on synchronized text and image interfaces in the context of multimodal human-AI collaboration workspaces.
Figure 1: The Pairit platform. On the left is the task panel, and on the right is the chat panel. In the Human-Human condition, chat messages and edits on the task panel, including text edits, image selections, and AI image generations, are synchronized in real-time. In the Human-AI condition, the participant chats with an AI agent with full context of the user interface (UI; see Section 2.2), and the AI can edit text, select images, and generate AI images.
2.2 AI agent
Context.
The platform uses OpenAI’s multimodal gpt-4o model (specifically gpt-4o-2024-08-06), which processes text and images in a single model. To give the AI full context of the task and collaboration, each API call to gpt-4o is prompted with the information on the screen so it has the full context of the collaboration. In the prompt, we include the same text of the task given to participants, previous submissions, current copy, elapsed time, history of its own actions, history of its chain-of-thought, chat history, and general instructions. Moreover, a screenshot is taken of the image after every change and included as an input so the AI can observe the image and its evolution over time and edits. These details enable human collaboration with a fully functional, multimodal AI agent.
Actions.
To ensure that the human-human and human-AI conditions are comparable, the AI agent can take the same actions a human participant except for submitting ads. The actions include sending messages, editing any element of the ad copy (including the headline, primary text, and description), selecting images, generating images using the Dall-E 3 interface, and waiting (i.e., not taking any action). The agent is prompted every 10 seconds whether to engage in action.
Chain-of-thought.
To ensure the agent is taking actions appropriately, we use chain-of-thought prompting (Wei et al., 2023). [^2] We used OpenAI’s structured output feature for chain-of-thought prompting. See their documentation. Specifically, we prompt the model with questions to reflect on the state of the collaboration. These questions were necessary to avoid undesirable model behavior, such as repeatedly sending the same message and were developed and perfected through trial and error.
@@ 2.3 Procedure
As soon as a participant was redirected to our platform from Prolific, they were randomized into either the human-human or the human-AI condition (Figure 2A). In the human-human condition, participants joined a queue until another participant was available, at which point they were paired with each other. In the human-AI condition, a participant joined a simulated queue, in which they waited for a random amount of time between 1 and 5 seconds, after which they were paired with an AI agent. We do not reveal whether or not the partner is a human or an AI until the post-task survey.
Figure 2: Overview of methods. (A) Participants are randomized into collaborating with another participant or an AI agent. (B) Participants collaborate with another participant or an AI agent to produce ads in a real-time collaborative workspace.
@@
After participants were paired with each other, they answered a 10-item survey, on a 7-point Likert scale, to measure their Big Five personality traits (Rammstedt and John, 2007).
@@
Once the participants were paired and completed a pre-task survey, they entered the collaborative workspace in which they could message each other and edit and submit ads (Figure 2B). The edits were synchronized and messages were transmitted between the participants in real-time using websockets (i.e., tiptap.dev and pusher.com), as they are in commercial collaboration tools (e.g., Google Docs) and chat applications (e.g., Slack). The participants had 40 minutes to submit as many ads as they could produce. At the end of the 40 minutes, the participants were automatically redirected to the post-task survey.
@@
After the completion of the ad creation task, the participants answered a 35-item teamwork quality survey (Hoegl and Gemuenden, 2001). The original survey consisted of 38 items, each on a 7-point Likert scale; however, we removed 3 items from the communication facet that did not apply to a single-session online collaboration task. In addition, we asked four questions regarding participants’ perception of AI, all on a 7-point Likert scale. These included two questions about their experience using AI (i.e., “I have used artificial intelligence (AI) chatbots before (e.g., ChatGPT, Bard).” and “I had a positive experience using AI chatbots.”), one question about their perception of their partner as an AI (i.e., “I believe my partner was an AI during the task.”) and a final question in which we revealed the identity of their partner (human or AI) and ask whether that changed the perception of the quality of the collaboration (i.e., “Your partner was [an AI assistant/a human]. Knowing this, to what extent has your perception of the quality of your collaboration changed?”). 85% of participants completed the post-task survey.